OpenStudy (anonymous):

Can someone help!!!?? The stem-and-leaf plot shows the heights in centimeters of Teddy Bear sunflowers grown in two different types of soil. Soil A Soil B 5 | 9 5 2 1 1| 6 | 39 510 | 7 | 023678 21| 8 | 3 0| 9 Key: 9|6 means 69,Key: 5|8 means 58 a) Describe the overlap of the two sets of data. b) Calculate the mean of each data set. c) Calculate the mean absolute deviation (MAD) of each data set. d) Which set is more variable? How do you know?

OpenStudy (goldphenoix):

I do not understand about problem a. For B just add up all of the number, and divide it by the amount that you added. So do 61 + 61 + 62 + 65... and so on... And make sure to do it for Soil B as well. C is a little bit tricky. (More like more work to do) Though C for Soil A is 8.16, and for Soil B is 5.4. If you want me to explain it, then tell me. For d, what does it mean by more variable?

OpenStudy (goldphenoix):

If you want me to do the mean for you, or explain the mean better, then let me know! :)

OpenStudy (goldphenoix):

For mean, MAD, both?

OpenStudy (goldphenoix):

Oh, to find MAD, you can use a few methods. So all of the methods take a while. I'm going to teach you this method. Step one: Calculate the mean of one of the soil group. Step two: Determine deviation of each variable from the Mean. So do: 61 - (The mean) = (Something) (1) 61 - (The mean) = (Something) (2) 62 - (The mean) = (Something) (3) And so on all the way to 90. Step three: Make the deviation 'absolute' by squaring and determining the roots i.e eliminate the negative aspect. So what you have to do is adding up all of the something you got and divide it by the amount of something. (I believe the amount of data in the set is 10. So after adding up all of the number, divide it by 10.) Step four: Do the WHOLE previous step on Soil B. (Seems like it can take another hour) Now this is one way you can find MAD. But this method is a little bit easier than the other.

OpenStudy (goldphenoix):

Do you understand what I did?^

OpenStudy (goldphenoix):

That's good. :)

OpenStudy (goldphenoix):

MAD takes forever, right?

OpenStudy (goldphenoix):

Deviation just means how far from the normal.

OpenStudy (goldphenoix):

By the way, for step 2, when the (something) is a negative, change it to positive. You're suppose to find the absolute value of all of the (something) Therefore, all of the (something) will become positive. So -10 will become 10. 20 will become 20. I hoe this help!